OpenStudy (anonymous):

Express with positive exponents. x ^{-1}y ^{-2}/z ^{-3}

OpenStudy (anonymous):

\[\frac{z^3}{x y^2} \]

OpenStudy (anonymous):

x ^{-1}y ^{-2}/z ^{-3} (1/x)(1/y^2)/(1/z^3) z^3/xy^2

OpenStudy (anonymous):

is it \[ \dfrac{x ^{-1}y ^{-2}}{z ^{-3}} \]

OpenStudy (anonymous):

If that is the question, then it is \[\dfrac{z^3}{xy^2}\]